OpenStudy (hackergod):

Math help? YOU WILL GET A MEDAL!!!!!(they are radical not square root)on all these problems if i say square root it could be radical, so if it dont make since, find the problem with radical rather than square root. (btw all the following signs should be the same )Thanks guys!! 3. Which of the following is equal to square root 50a^6b^7 A. 5a^3b^3 square root 2b B. 5a^2b^3 square root 2b C. 10a^3b^3 D. 5a^3b^3

OpenStudy (hackergod):

4. What is the simplist form of the expression? Squareroot3-squareroot6/squareroot3+squ... A.-1-2squareroot18/3 b. -3-2squareroot18/9 c.-3+2squareroot2 d. 9-2 squareroot 18

OpenStudy (hackergod):

7. what is the solution to x=squareroot2x+3 A.3 B.-1 C. 3 and -1 D -3 and 1 9. graph y=squareroot 3x-5-1

OpenStudy (hackergod):

11. what is the inverse of f if f(x)= ^3 squareroot x-5? Please can you show all steps for these. 16. Multiply and simplify if possible. Radical 7(radical x – 7 radical 7) 17. What is the solution of the equation? 3^5radical (x+2)^3+3=27

OpenStudy (hackergod):

18. What is the solution of the equation? Radical 2x+13-5=x 19. Divide and simplify id possible. Radical 250x^16 / radical 2x 20. What is the simplest form of the radical expression? 3^3radical2a-6^3radical2a 21. Simplify. 3^1/2 *3^1/2 22. Rewrite ^nradical x^m using a rational expression

OpenStudy (hackergod):

help!!!???

OpenStudy (austinl):

Okay, first off... try and only ask a couple of questions at once. Then later ask if you can ask some more. Lets start out with the first one shall we?

OpenStudy (hackergod):

okay, sorry, and thanks for helping

OpenStudy (austinl):

\(\Large{\sqrt{50a^6b^7}}\) This is what the first one looks like right?

OpenStudy (hackergod):

yes

OpenStudy (hackergod):

A. 5a^3b^3 square root 2b B. 5a^2b^3 square root 2b C. 10a^3b^3 D. 5a^3b^3 are the options

OpenStudy (austinl):

I am not getting any of those.

OpenStudy (hackergod):

its not square root its radical for all of the problems..

OpenStudy (austinl):

Square Root = Radical

OpenStudy (hackergod):

ohh.... then we can move on??

OpenStudy (hackergod):

4. What is the simplist form of the expression? Squareroot3-squareroot6/squareroot3+squ... A.-1-2squareroot18/3 b. -3-2squareroot18/9 c.-3+2squareroot2 d. 9-2 squareroot 18

OpenStudy (austinl):

What is the end of that top part?

OpenStudy (hackergod):

+squareroot 6

OpenStudy (austinl):

\(\displaystyle \frac{\sqrt{3}-\sqrt{6}}{\sqrt{3}+\sqrt{6}}\) Is this correct?

OpenStudy (hackergod):

Yes(:

OpenStudy (austinl):

I am not quite sure how to explain this.... Try C, see if that works.

OpenStudy (hackergod):

I can't " Try" it... its somthing i have tosubmit

OpenStudy (hackergod):

@jjuden

OpenStudy (jjuden):

do you need help with all of them or just one??

OpenStudy (hackergod):

alll...... thank you so much!!

OpenStudy (jjuden):

ok hold on

OpenStudy (hackergod):

ok thanks

OpenStudy (jjuden):

ok i tried to do them but the computer told me i had mine wrong and wont tell me wats right sorry i cant help you this is the area that im worst at so sorry ):

OpenStudy (hackergod):

do you know anyne who is good with this???

OpenStudy (jjuden):

my brother buts hes being an retriceole right now sorry

OpenStudy (jjuden):

lol a$$hole not wat the website corrected

OpenStudy (hackergod):

tell him he would really be helping someone.....

OpenStudy (hackergod):

does he have an account?

OpenStudy (hackergod):

@Gin_Ichimaru

OpenStudy (hackergod):

@phi

OpenStudy (hackergod):

@undeadknight26

OpenStudy (hackergod):

@austinL !!! plzz...

OpenStudy (hackergod):

@athenasannabeth

OpenStudy (hackergod):

@amistre64

OpenStudy (hackergod):

@ranga

undeadknight26 (undeadknight26):

thats alot of questions...sorry i do not have time right now for so many...

OpenStudy (hackergod):

can you just do a few then? how ever many you can...

OpenStudy (hackergod):

@undeadknight26 do whatever yppu can plz

OpenStudy (hackergod):

@twerk ...

OpenStudy (hackergod):

@ganeshie8 help??

OpenStudy (hackergod):

@Preetha @.Sam. @phi @Zarkon @ash2326 @hartnn PLEASE HELP!! GGET MEDAL!

OpenStudy (hackergod):

@MrNebula

OpenStudy (hackergod):

@mukushla @mathstudent55

OpenStudy (hackergod):

@mathman806 can you ?

OpenStudy (studygurl14):

@harckergod that is easy. All you do is plug in numbers for a and b to see which answer choices are equal to the original one and which isn't

OpenStudy (studygurl14):

@hackergod

OpenStudy (hackergod):

for all ??

OpenStudy (studygurl14):

For the first one.

OpenStudy (hackergod):

.... ok soo what would it be? B?

OpenStudy (hackergod):

helppppppppppppppppp!!! Please

OpenStudy (hackergod):

@amistre64

OpenStudy (hackergod):

just ansewr ones you know or can do... @amistre64 @rocker128

OpenStudy (anonymous):

is #11